OpenStudy (blackops2luvr):

A cone-shaped paper cup can hold 35 cubic inches of food. If the radius of the cone is 2.4 inches, what is the height of the cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 5.81 inches 6.57 inches 7.89 inches 8.89 inches @iGreen @AnswerMyQuestions

OpenStudy (anonymous):

1/3 x 3.14 x 2.4. Then divide 35 by that.

OpenStudy (igreen):

1/3 × pi × r^2 × h 35 = 1/3 × 3.14 × 2.4^2 × h 35 = 1/3 × 3.14 × 5.76 × h 35 = 6.0288h 35 / 6.0288 = ?

OpenStudy (igreen):

@blackops2luvr Divide that.

OpenStudy (anonymous):

@blackops2luvr 1/3 x 3.14 x 2.4. Then divide 35 by that.

OpenStudy (blackops2luvr):

5.8054670913

OpenStudy (anonymous):

Yeah

OpenStudy (igreen):

Yep

OpenStudy (anonymous):

WAIT

OpenStudy (blackops2luvr):

what i got 100pc

OpenStudy (igreen):

That's right.

OpenStudy (blackops2luvr):

igreen dont give the medal to me

OpenStudy (igreen):

Why not?

OpenStudy (anonymous):

Oh.

OpenStudy (blackops2luvr):

give it to @AnswerMyQuestions

OpenStudy (igreen):

No, I won't.

OpenStudy (anonymous):

:O

OpenStudy (blackops2luvr):

He deserves it more than me :)

OpenStudy (igreen):

He should arrest himself because he's a cheater..

OpenStudy (anonymous):

How?

OpenStudy (igreen):

Posting your FLVS questions here and giving direct answers.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

When was the last time I posted an FLVS question? And gave a direct answer?

OpenStudy (igreen):

@AnswerMyQuestions was wrong.. he said to multiply 1/3 x 3.14 x 2.4 and divide by 35. 1/3 x 3.14 x 2.4 = 2.512 2.512 / 35 = 0.07 @blackops2luvr

OpenStudy (blackops2luvr):

GUYS STOP ARGUING MY GOD

OpenStudy (igreen):

Plus it was supposed to be 1/3 x 3.14 x 2.4^2

OpenStudy (igreen):

So why did you give him a medal? Plus I actually EXPLAINED THE STEPS.

OpenStudy (anonymous):

Ok, you don't have to get mad when you don't get the medal little baby.

OpenStudy (anonymous):

It was 1 mistake.

OpenStudy (igreen):

Yeah, it could've gave him the wrong answer.

OpenStudy (anonymous):

As iGreen would say: ARGH! STOP STEALING MY QUESTIONS!

OpenStudy (igreen):

Ah, you can smell the rage in the air. :)

OpenStudy (blackops2luvr):

not you iGreen

OpenStudy (igreen):

...................

OpenStudy (igreen):

LOLOLOLOL

OpenStudy (anonymous):

I don't even care that much about medals. You go crazy for them... And the rage when he doesn't get the medal... Wow...

OpenStudy (blackops2luvr):

Brotherly love :')

OpenStudy (igreen):

Heh. :P

OpenStudy (piercetheveil47):

was A the answer? lol